# Tessellated phase diagram and extended criticality in driven quasicrystals and quantum Hall matter

### v2

@@ Abstract
The well-known mapping between 1D quasiperiodic systems and 2D integer quantum Hall matter can also be applied in the presence of driving. Here we explore the effect of time-varying electric fields on the transport properties and phase diagram of Harper-Hofstadter materials. We consider light of arbitrary polarization illuminating a 2D electron gas at high magnetic field; this system maps to a 1D quasicrystal subjected to simultaneous phasonic and dipolar driving. We show that this generalized driving generates a tessellated phase diagram featuring a nested duality-protected pattern of metal-insulator transitions. Circularly or elliptically polarized light can create an extended critical phase, opening up a new route to achieving wavefunction multifractality without fine-tuning to a critical point, as well as induce Floquet topological insulators. We describe in detail a path to experimental realization of these phenomena using lattice-trapped ultracold atoms.
@@
Quasiperiodic systems host a rich array of physical phenomena ranging from localization [1, 2, 3, 4] to fractal spectra [5, 6, 2] to criticality [7, 8, 9, 10, 11] to non-trivial topology [12, 13, 14, 15, 16]. Still richer possibilities arise in the presence of external driving, including a competition between thermalization and many-body localization [17, 18] and potentially extended critical phases featuring multifractal wavefunctions [19, 20, 21, 18, 22, 23, 24, 25, 26, 27, 27, 11, 28, 29, 30]. These phenomena arise from the long-range spatial correlation of quasiperiodicity, which itself is the result of projecting a higher-dimensional periodic structure into the lower-dimensional physical space  [31, 12, 13, 14, 15, 16]. The connection to the higher-dimensional space also gives rise to a degree of freedom called a phason mode. A recent experiment [32] illustrated a connection between two seemingly unrelated localization phenomena by mapping a rapidly oscillating phasonic modulation to linearly polarized monochromatic light in the higher-dimensional space. It is then natural to investigate phenomena induced by arbitrary polarizations of irradiation in the higher-dimensional space.
@@
In this work, we complete this picture by considering simultaneous dipolar [33] and phasonic [34, 32] modulation in the physical space. The phase difference between these two modulations maps to the polarization of light illuminating the higher-dimensional space, potentially enabling quantum simulation in a one-dimensional system of the effects of arbitrarily-polarized optical driving of a 2D electron gas at high magnetic field. We report the following main results. First, the combination of phasonic and dipolar modulation allows coherent control of both the tunneling [33] and quasi-disorder strengths [32], resulting in a tessellated localization phase diagram with interlaced localized and delocalized phases. We note that this coherent control toolset allows all elements of the non-interacting Hamiltonian to be flipped in sign, which would effectively reverse the direction of the flow of time. Further, by considering circularly polarized illumination in the higher-dimensional space, we demonstrate that an extended exotic phase hosting multifractal wavefunctions emerges naturally and without fine-tuning [25]. Additionally, we show that for commensurate (periodic) cases, such driving protocols lead to topological edge modes. We identify a interesting connection between the generalized Rice-Mele and Su-Schrieffer-Heeger models; and the driven Harper-Hofstadter model. Finally, we propose a straightforward experimental realization of all these phenomena using ultracold atoms in a driven bichromatic optical lattice.
@@
This paper is organized as follows. In section I we review the projection of quasiperiodic systems from higher-dimensional space. Readers familiar with the mapping can skip to Section I.1 where we extend this mapping to include spatially homogeneous irradiation in the higher-dimensional space and comment on the effect of its polarization in the 1D space. We discuss the drive-induced tessellated localization phase diagram in Section II.1, the effective time-reversal protocol in Section II.2, Floquet-engineering of critical phases in Section II.3, and topological edge modes in Section III. The experimental realization with driven ultracold atoms is discussed in Section IV. We conclude by noting some intriguing future directions in Section V.
@@ I Quasiperiodicity from higher dimensions
where $J$ is the tunneling strength, $\Delta$ is the strength of the quasiperiodic potential, $\hat{c}^{(\dagger)}_{j}$ is the spinless annihilation (creation) operator at the site $j$, $\hat{n}_{j}:=\hat{c}_{j}^{\dagger}\hat{c}_{j}$, and $\beta$ is the incommensurate ratio that defines the quasiperiodicity, i.e. the ratio of the spatial period of the underlying lattice to that of the quasiperiodic modulation. The AAH model supports a localization phase transition at $|\Delta/J|=2$ when $\beta$ is irrational [1, 3, 4]. When $\Delta>2J$, all eigenstates are localized and the system is in the localized phase; when $\Delta<2J$, all eigenstates are extended and the system is in the delocalized phase. At the critical point, all eigenstates are multifractal with non-trivial scaling properties.
@@
The incommensurate ratio $\beta$ then represents the ratio of the magnetic flux over the flux quantum in the 2D picture, and the quasiperiodic strength $\Delta$ becomes the tunneling along the extra dimension. Historically, this mapping was first recognized by Harper leading to the Harper equation [31] and later the Hofstadter butterfly [5].
Figure 2: The quasienergy diagrams within the central Floquet-Brillouin zone of the DDAAH model under different polarizations: (a) left-handed circular polarization (LCP); (b): linearly-polarized and (c) right-handed circular polarization (RCP). The color represents the edge-locality marker $\mathcal{B}$ (see Eq.(9)) that identifies the edge states in the spectra localized at the right (R) or left (L) end. Here $\Delta/J=2$, $K_{0}=\varphi_{0}=1.5$, $\hbar\omega=20J$, $\kappa=0$ and $\theta_{x}=\pi/2$ with 100 lattice sites and 5 Floquet-Brillouin zones.
@@ I.1 From irradiated Harper-Hofstadter Lattice to Doubly Driven AAH Quasicrystal
    When $\delta\theta=\pm\pi/2$, the irradiation is elliptically polarized. The semi-major and -minor axes of the polarization ellipse are parallel to the lattice’s axes. Additionally, the irradiation becomes circularly polarized when $K_{0}=\varphi_{0}$.
@@
We see that dimensional reduction has mapped the field component parallel to the physical dimension (labeled by $j$) into a real oscillating field in 1D with amplitude $K$. We refer to this modulation as the dipolar modulation, since it has the form of an electric dipole force. The field along the extra dimension (labelled by $l$) is projected into a phasonic modulation [32] with amplitude $\varphi_{0}$. This correspondence of parameters between 1D and 2D physical models is schematically shown in Fig 1. We can thus interpret a time-dependent phason as a virtual electric field along the extra dimension [37, 38, 39, 14, 42, 32]. This interpretation underpins, for example, topological pumping in the bichromatic lattice [12, 42]: adiabatic linear scanning of the phason maps to a static electric field (and thus charge-pumping) along the extra dimension in the higher-dimensional HH model. For our purposes, a particularly notable feature of this mapping is that the polarization of the radiation in the 2D space maps to the phase difference between the dipolar and phasonic modulations in 1D. In light of this, we will use the terminology of polarization to refer to these modulations for the remainder of the paper.
@@
As Fig. 2 demonstrates, the impact of different polarizations is evident in the quasienergy spectra as a function of $\beta$ (the Floquet-Hofstadter butterfly) [40, 43, 44]. We observe that linearly-polarized modulation preserves the structure of the Hofstadter butterfly (Fig. 2(b)), while circularly-polarized modulation asymmetrically distorts it (Fig. 2(a) and (c)). This is due to the different symmetry-breaking properties of the incident light in the 2D space. Circularly polarized light breaks both time-reversal and sublattice symmetries in the higher-dimensional space. Consequently, the corresponding symmetries in the Hofstadter butterfly spectrum (reflection symmetry about $\beta=1/2$ and particle-hole symmetry about $E=0$, respectively) are broken  [40]. On the other hand, linearly polarized light does not break these symmetries [45, 40]. Ref. [40] also noted other small differences of the Floquet-Hofstadter butterfly spectra compared to the time-averaged prediction. These differences, however, are not due to symmetry-breaking from the polarization, but deviations from the time-averaged description, which only strictly holds at $\omega\rightarrow\infty$.
@@
Much of the physics of the 1D model (Eq. (1)) can be understood in the 2D space picture, even without any explicit calculations. First, time-periodic driving modifies the localization properties of the system. It is well-known that a lattice system under time-periodic driving exhibits dynamic localization [45, 33]. This phenomenon can be intuitively understood as time-averaged Bloch oscillation in an AC electric field. Depending on the field strength projected to each dimension, the driving coherently re-scales the tunneling strength along that dimension, and can stop the tunneling of the particle along that dimension when the driving amplitude equals certain resonant values. This competition of tunneling suppression between different axes in the higher-dimensional space will impact the localization properties in the projected physical space, as the tunneling strength along the extra dimension becomes the strength of the quasiperiodic chemical potential variation through dimensional reduction. As we will see in Sections II.1 and II.2, this mechanism leads to a rich tessellated localization phase diagram in the space of drive polarizations, and allows for coherent and time-reversible control of localization dynamics.
@@
Circularly-polarized light will in general lead to next-nearest-neighbor (NNN) hopping in the higher-dimensional space through virtual photon absorption and emission processes [46]. When projected to the 1D physical space, the NNN hopping can lead to mobility edges and critical states [8, 14]. We will show in Section II.3 that the NNN hopping in the 2D higher-dimensional space allows Floquet engineering of an extended critical phase hosting multifractal wave functions without fine-tuning.
Finally, external periodic driving can modify the topology of the system, leading to Floquet topological insulators [47]. Given that our primary focus here is the various localization phenomena in the driven quasiperiodic model, we will provide just a brief discussion of topology in section III, where we connect the driven Harper-Hofstadter model to the generalized Rice-Mele and Su–Schrieffer–Heeger models.
@@ II Floquet-Engineering of Localization Transition and Critical Phases
Figure 4: Effective time-reversal dynamics of the DDAAH model. The system is driven for 1000 cycles for forward propagation at $K_{0}^{\mathrm{for}}\approx 1.6965$ ($\mathcal{J}_{0}(K_{0}^{\mathrm{for}})=0.4$) and another 1000 cycles for backward propagation at $K_{0}^{\mathrm{back}}\approx 3.7152$ ($\mathcal{J}_{0}(K_{0}^{\mathrm{back}})=-0.4$), shown as the red dots in the inset of (a). The time-reversal dynamics are observed in both (a) the time-evolution of the density profile and (b) the return probability $\mathcal{L}(t)$. The initial state is a single-site excitation $\psi(t)=\delta_{j,0}$, with parameters $\Delta/J=1$, $\theta_{x}=\theta_{y}=0$ and $\hbar\omega=100$. The Hamiltonian (Eq.2) is numerically integrated with the second order split-operator method with a time step of $dt\approx 1.26\times 10^{-5}\hbar/J$.
@@ II.1 Tessellated phase diagram
where $L$ is the system size and $\psi_{j}^{(i)}$ is the $i^{\mathrm{th}}$ Floquet eigenstate in the position basis. A (non)vanishing $\mathrm{IPR}_{2}^{(i)}$ indicates a spatially extended (localized) state. We further define $\langle\mathrm{IPR}_{2}\rangle$, obtained by averaging the $\mathrm{IPR}_{2}^{(i)}$ over all Floquet eigenstates in the central Floquet-Brillouin zone which is defined by $-\hbar\omega/2\leq\varepsilon\leq\hbar\omega/2$. The results are shown in Fig. 3. The phase boundaries predicted by the analytical high-frequency approximation (Eq. 4) are shown as dashed curves and demonstrate excellent agreement with numerical results. The phase diagram has a tessellated self-dual structure with interlacing areas of localized and delocalized phases due to the competition between $\Delta_{\mathrm{eff}}$ and $J_{\mathrm{eff}}$. We emphasize that this rich localization phase diagram characterizes, and could be observed in, both a 1D quasicrystal subjected to combined dipolar and phasonic modulation, and a 2D electron gas at high magnetic field and strong irradiation.
@@ II.2 Effective time-reversal dynamics
These results demonstrate that the localization quantum phase transition of the 1D DDAAH model can be coherently controlled by combined dipolar and phasonic modulations. In contrast to previous schemes where either the dipolar or the phasonic modulation is individually applied [33, 32], the use of combined modulations allows for simultaneous control of the tunneling strength and quasiperiodic potential together, opening up new possibilities.
As an example, we describe a driving protocol that achieves effective time-reversal dynamics in the DDAAH model. For simplicity, here we consider both modulations to have the same amplitude ($\varphi_{0}=K_{0}$). By simply choosing a sequence of two modulation amplitudes arranged around the zeros of the Bessel function $\mathcal{J}_{0}(K_{0})$ such that $\mathcal{J}_{0}(K_{0}^{\mathrm{for}})=-\mathcal{J}_{0}(K_{0}^{\mathrm{back}})$, the signs of all parameters in the effective Hamiltonian $\hat{H}^{(0)}$ are flipped, effectively reversing the direction of the flow of time. Results of numerical integration of the discrete time-dependent Schrödinger equation of the Hamiltonian (2) are shown in figure 4. The time-reversal is evident from both the wave function propagation and the return probability, defined as $\mathcal{L}(t):=|\langle\psi(t)|\psi(0)\rangle|^{2}$. Such techniques may open up the possibility to experimentally study dynamical quantum phase transitions in Anderson-localized matter via measurements of the Loschmidt echo [54].

### v3

# Tunably-polarized driving light controls the phase diagram of 1D quasicrystals and 2D quantum Hall matter

@@ Abstract
The well-known mapping between 1D quasiperiodic systems and 2D integer quantum Hall matter can also be applied in the presence of driving. Here we explore the effect of time-varying electric fields on the transport properties and phase diagram of Harper-Hofstadter materials. We consider light of arbitrary polarization illuminating a 2D electron gas at high magnetic field; this system maps to a 1D quasicrystal subjected to simultaneous phasonic and dipolar driving. We show that this generalized driving generates a tessellated phase diagram featuring a nested duality-protected pattern of metal-insulator transitions. Circularly or elliptically polarized light can create an extended critical phase, opening up a new route to achieving wavefunction multifractality without fine-tuning to a critical point. We describe in detail a path to experimental realization of these phenomena using lattice-trapped ultracold atoms.
Quasiperiodic systems host a rich array of physical phenomena ranging from localization [1, 2, 3, 4] to fractal spectra [5, 6, 2] to criticality [7, 8, 9, 10, 11] to non-trivial topology [12, 13, 14, 15, 16]. Still richer possibilities arise in the presence of external driving, including a competition between thermalization and many-body localization [17, 18] and potentially extended critical phases featuring multifractal wavefunctions [19, 20, 21, 18, 22, 23, 24, 25, 26, 27, 27, 11, 28, 29, 30]. Identifying a simple experimental control knob that tunes and connects these dynamical phenomena is the open challenge we address in this theoretical manuscript.
@@
All these phenomena arise from the long-range spatial correlation of quasiperiodicity, which itself is the result of projecting a higher-dimensional periodic structure into a lower-dimensional physical space  [31, 12, 13, 14, 15, 16]. The connection to the higher-dimensional space also gives rise to a degree of freedom called a phason mode. A recent experiment [32] illustrated a connection between two apparently unrelated localization phenomena — dynamical and Aubry-André localization — by mapping a rapidly oscillating phasonic modulation to linearly polarized monochromatic light in the higher-dimensional space. This naturally raises the question of what the effects are of arbitrarily tuning the polarization of irradiation in the higher-dimensional space.
@@
In this work, we theoretically demonstrate that tunably polarized driving can be used to control localization, generate extended critical phases, and induce Floquet topological insulators by tuning the system along the axes of an intricately tesselated phase diagram in the space of drive polarizations. Specifically, we consider simultaneous dipolar [33] and phasonic [34, 32] modulation in the 1D physical space. The phase difference between these two modulations maps to the polarization of applied light in the higher-dimensional space, enabling quantum simulation in a one-dimensional system of the effects of arbitrarily-polarized optical driving of a 2D electron gas at high magnetic field. We report the following main results. First, the combination of phasonic and dipolar modulation allows coherent control of both the tunneling [33] and quasi-disorder strengths [32], resulting in a tessellated localization phase diagram with interlaced localized and delocalized phases. We note that this control toolset allows all elements of the non-interacting Hamiltonian to be flipped in sign, effectively reversing the direction of the flow of time. We demonstrate that circularly polarized illumination naturally and without fine-tuning generates an extended exotic phase hosting multifractal wavefunctions [25]. Distinct from previous work on drive-induced extended criticality [22, 35, 36], this result does not rely on mixing localized and ballistic modes, but is a direct consequence of the drive polarization. Finally, we propose a straightforward experimental realization of all these phenomena using ultracold atoms in a driven bichromatic optical lattice.
@@
The paper is organized as follows. In Section I we review the projection of quasiperiodic systems from higher-dimensional space. Readers familiar with the mapping can skip to Section I.1 where we extend this mapping to include spatially homogeneous irradiation in the higher-dimensional space and comment on the effect of its polarization in the 1D space. We discuss the drive-induced tessellated localization phase diagram in Section II.1, the effective time-reversal protocol in Section II.2, and Floquet-engineering of critical phases in Section II.3. A possible experimental realization with driven ultracold atoms is discussed in Section III. We conclude by noting some intriguing future possible directions in Section IV.
@@ I Quasiperiodicity from higher dimensions
where $J$ is the tunneling strength, $\Delta$ is the strength of the quasiperiodic potential, $\hat{c}^{(\dagger)}_{j}$ is the spinless annihilation (creation) operator at the site $j$, $\hat{n}_{j}:=\hat{c}_{j}^{\dagger}\hat{c}_{j}$, and $\beta$ is the incommensurate ratio that defines the quasiperiodicity, i.e. the ratio of the spatial period of the underlying lattice to that of the quasiperiodic modulation.
The AAH model supports a localization phase transition at $|\Delta/J|=2$ when $\beta$ is irrational [1, 3, 4]. When $\Delta>2J$, all eigenstates are localized and the system is in the localized phase; when $\Delta<2J$, all eigenstates are extended and the system is in the delocalized phase. At the critical point, all eigenstates are multifractal with non-trivial scaling properties.
@@
The incommensurate ratio $\beta$ then represents the ratio of the magnetic flux over the flux quantum in the 2D picture, and the quasiperiodic strength $\Delta$ becomes the tunneling along the extra dimension. Historically, this mapping was first recognized by Harper [31] and later explored in the context of the famous Hofstadter butterfly energy spectrum [5].
Figure 2: Quasienergy spectra of the DDAAH model for different drive polarizations: (a) left-handed circular polarization (LCP); (b): linear polarization; and (c) right-handed circular polarization (RCP). The color represents the edge-locality marker $\mathcal{B}$ (see Eq.(13)) that identifies the edge states in the spectra localized at the right (R) or left (L) end. Here $\Delta/J=2$, $K_{0}=\varphi_{0}=1.5$, $\hbar\omega=20J$, $\kappa=0$ and $\theta_{x}=\pi/2$ with 100 lattice sites and 5 Floquet-Brillouin zones.
@@ I.1 From irradiated Harper-Hofstadter Lattice to Doubly Driven AAH Quasicrystal
    When $\delta\theta=\pm\pi/2$, the irradiation is elliptically polarized. The semi-major and -minor axes of the polarization ellipse are parallel to the lattice axes. Additionally, the irradiation becomes circularly polarized when $K_{0}=\varphi_{0}$.
@@
We see that dimensional reduction has mapped the field component parallel to the physical dimension (labeled by $j$) into a real oscillating field in 1D with amplitude $K$. We refer to this modulation as the dipolar modulation, since it has the form of an electric dipole force. The field along the extra dimension (labelled by $l$) is projected into a phasonic modulation [32] with amplitude $\varphi_{0}$. This correspondence of parameters between 1D and 2D physical models is schematically shown in Fig 1. We can thus interpret a time-dependent phason as a virtual electric field along the extra dimension [39, 40, 41, 14, 44, 32]. This interpretation underpins, for example, topological pumping in the bichromatic lattice [12, 44]: adiabatic linear scanning of the phason maps to a static electric field (and thus charge-pumping) along the extra dimension in the higher-dimensional HH model. For our purposes, a crucial feature of this mapping is that the polarization of the radiation in the 2D space maps to the phase difference between the dipolar and phasonic modulations in 1D. In light of this, we will use the terminology of polarization to refer to these modulations for the remainder of the paper.
@@
As Fig. 2 demonstrates, the impact of different polarizations is evident in the quasienergy spectra as a function of $\beta$ (the Floquet-Hofstadter butterfly) [42, 45, 46]. We observe that linearly-polarized modulation preserves the structure of the Hofstadter butterfly (Fig. 2(b)), while circularly-polarized modulation asymmetrically distorts it (Fig. 2(a) and (c)). This is due to the different symmetry-breaking properties of these different polarizations of the incident light in the 2D space. Circularly polarized light breaks both time-reversal and sublattice symmetries in the higher-dimensional space; consequently, the corresponding symmetries in the Hofstadter butterfly spectrum (reflection symmetry about $\beta=1/2$ and particle-hole symmetry about $E=0$, respectively) are broken  [42]. On the other hand, linearly polarized light does not break these symmetries [47, 42]. Ref. [42] also noted other small differences of the Floquet-Hofstadter butterfly spectra compared to the time-averaged prediction. These differences, however, are not due to symmetry-breaking from the polarization, but deviations from the time-averaged description, which only strictly holds at $\omega\rightarrow\infty$.
@@
Much of the physics of the 1D model (Eq. (1)) can be understood in the 2D picture, even without any explicit calculation. First, time-periodic driving modifies the localization properties of the system. It is well-known that a lattice system under time-periodic driving exhibits dynamic localization [47, 33]. This phenomenon can be intuitively understood as time-averaged Bloch oscillation in an AC electric field. Depending on the field strength projected to each dimension, the driving coherently re-scales the tunneling strength along that dimension, and can stop the tunneling of the particle along that dimension when the driving amplitude equals certain resonant values. This competition of tunneling suppression between different axes in the higher-dimensional space will impact the localization properties in the projected physical space, as the tunneling strength along the extra dimension becomes the strength of the quasiperiodic chemical potential variation through dimensional reduction. As we will see in Sections II.1 and II.2, this mechanism leads to a rich tessellated localization phase diagram in the space of drive polarizations, and allows for coherent and time-reversible control of localization dynamics.
@@
Moreover, driving with circularly-polarized light will in general lead to effective next-nearest-neighbor (NNN) hopping in the higher-dimensional space through virtual photon absorption and emission processes [48]. When projected to the 1D physical space, the NNN hopping can lead to mobility edges and critical states [8, 14]. We show in Section II.3 that the NNN hopping in the 2D higher-dimensional space allows Floquet engineering of an extended critical phase hosting multifractal wave functions without fine-tuning. This polarization-induced effective NNN hopping also modifies the topology of the system due to broken time-reversal symmetry. Given that our primary focus here is on localization phenomena in the driven quasiperiodic model, we will provide just a brief discussion of topology in Appendix D, where we connect the driven Harper-Hofstadter model to the generalized Rice-Mele and Su–Schrieffer–Heeger models.
@@ II Floquet-Engineering of Localization Transition and Critical Phases
Figure 4: Controlling effective time-reversal in the DDAAH model. The system is driven for 1000 cycles for forward propagation at $K_{0}^{\mathrm{for}}\approx 1.6965$ ($\mathcal{J}_{0}(K_{0}^{\mathrm{for}})=0.4$) and another 1000 cycles for backward propagation at $K_{0}^{\mathrm{back}}\approx 3.7152$ ($\mathcal{J}_{0}(K_{0}^{\mathrm{back}})=-0.4$), shown as the red dots in the inset of (a). The time-reversal dynamics are observed in both (a) the time-evolution of the density profile and (b) the return probability $\mathcal{L}(t)$. The initial state is a single-site excitation $\psi(t)=\delta_{j,0}$, with parameters $\Delta/J=1$, $\theta_{x}=\theta_{y}=0$ and $\hbar\omega=100$. The Hamiltonian (Eq.2) is numerically integrated with the second order split-operator method with a time step of $dt\approx 1.26\times 10^{-5}\hbar/J$.
@@ II.1 Tessellated phase diagram
where $L$ is the system size and $\psi_{j}^{(i)}$ is the $i^{\mathrm{th}}$ Floquet eigenstate in the position basis. A (non)vanishing $\mathrm{IPR}_{2}^{(i)}$ indicates a spatially extended (localized) state. We further define $\langle\mathrm{IPR}_{2}\rangle$, obtained by averaging the $\mathrm{IPR}_{2}^{(i)}$ over all Floquet eigenstates in the central Floquet-Brillouin zone which is defined by $-\hbar\omega/2\leq\varepsilon\leq\hbar\omega/2$. The results are shown in Fig. 3. The phase boundaries predicted by the analytical high-frequency approximation (Eq. 4) are shown as dashed curves and demonstrate excellent agreement with numerical results. The phase diagram has a tessellated self-dual structure with interlacing areas of localized and delocalized phases due to the competition between $\Delta_{\mathrm{eff}}$ and $J_{\mathrm{eff}}$. We emphasize that this rich localization phase diagram characterizes, and could be observed in, both a 1D quasicrystal subjected to combined dipolar and phasonic modulation, and a 2D electron gas at high magnetic field and strong irradiation [55].
@@ II.2 Effective time-reversal dynamics
These results demonstrate that the localization quantum phase transition of the 1D DDAAH model can be coherently controlled by combined dipolar and phasonic modulation. In contrast to previous schemes where either the dipolar or the phasonic modulation is individually applied [33, 32], the use of combined modulation allows for simultaneous control of the tunneling strength and quasiperiodic potential together, opening up new possibilities. As an example of such a possibility, we describe a driving protocol that achieves effective time-reversal dynamics in the DDAAH model. This is potentially of practical as well as intrinsic interest; the ability to effectively reverse time evolution is a key component of schemes for experimentally probing fidelity decay or scrambling through Loschmidt echo or out-of-time-order correlators [56].
For simplicity, we consider both modulations to have the same amplitude ($\varphi_{0}=K_{0}$). By simply choosing a sequence of two modulation amplitudes arranged around the zeros of the Bessel function $\mathcal{J}_{0}(K_{0})$ such that $\mathcal{J}_{0}(K_{0}^{\mathrm{for}})=-\mathcal{J}_{0}(K_{0}^{\mathrm{back}})$, the signs of all parameters in the effective Hamiltonian $\hat{H}^{(0)}$ are flipped, effectively reversing the direction of the flow of time. Results of numerical integration of the discrete time-dependent Schrödinger equation of the Hamiltonian (2) are shown in figure 4, for parameters and hold times well within the range of experimental feasibility. The time reversal is evident from both the wave function propagation and the return probability, defined as $\mathcal{L}(t):=|\langle\psi(t)|\psi(0)\rangle|^{2}$.